# On Selection of Cross-Section Averages in Non-stationary Environments

### v2

@@ 6.3 Part $M\subset M_{0}$
$\leq\underbrace{T^{-(1+\tau)/2}N^{-1/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{o_{p}(C_{N,T}^{-2})}$
@@
$+\underbrace{N^{-1/2}T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\sqrt{N}T^{-(1+\tau)/2}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(N^{-1/2}T^{-1})}$
@@
$+\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(C_{N,T}^{-2})}$
@@
$+\underbrace{T^{(\tau-1)/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(T^{\tau/2}C_{N,T^{\tau}}^{-2})}$
@@
×$\times\underbrace{\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(1)}$
$=O_{p}(C_{N,T}^{-2}),$ (A.20)
@@
$\left\|\mathbf{I}\right\|=\left\|\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right\|=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}),$ (A.21)
@@
$\left\|\mathbf{I}\mathbf{I}\right\|=\left\|\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right\|=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}),$ (A.22)
@@
$\mathbf{III}$ $=\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})\mathbf{X}_{i}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}-\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}+\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}-\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{X}_{i}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
@@
where the term $\mathbf{III}_{d}$ is particularly problematic. By using random loadings assumption, it explicitly can be written as
$\mathbf{III}_{d}$ $=\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$=T^{-1}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}$
$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}$
$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}$
$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}$
$=\mathbf{III}^{1}_{d}+\mathbf{III}^{2}_{d}+\mathbf{III}_{d}^{3}+\mathbf{III}^{4}_{d}.$ (A.24)
Note that the key component in all $\mathbf{III}^{1}_{d}-\mathbf{III}^{4}_{d}$ is $\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$, where
@@
as $T\to\infty$, which is positive definite. Now, moving on to $\mathbf{III}^{1}_{d}$, we obtain
$\mathbf{III}^{1}_{d}$ $=T^{-1}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$
$=T^{\tau}\boldsymbol{\Gamma}^{\prime}\underbrace{(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}})}_{O_{p}(1)}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\underbrace{(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}})}_{O_{p}(1)}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}\underbrace{(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}})}_{O_{p}(1)}$
$=O_{p}(T^{\tau}),$ (A.26)
which means that the whole term is divergent, and the same applies to $\mathbf{III}^{1}_{d}$. In particular, $\mathbf{III}^{4}_{d}$ has the same order, since

### v3

@@ 5.3.1 Case of $M\subset M_{0}$
$\leq\underbrace{T^{-(1+3\tau)/2}N^{-1/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{O_{p}(T^{-(1+3\tau)/2}N^{-1/2}C_{N,T}^{-2})}$
@@
$+\underbrace{N^{-1/2}T^{-1-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\sqrt{N}T^{-(1+\tau)/2}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(N^{-1/2}T^{-1-\tau})}$
@@
$+\underbrace{T^{-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(T^{-\tau}C_{N,T}^{-2})}$
@@
$+\underbrace{T^{-(\tau+1)/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(T^{-(\tau+1)/2}\xi_{N,T,\tau}^{-1})}$
@@
$\times\underbrace{\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(1)}$
$=O_{p}(T^{-\tau}C_{N,T}^{-2})+O_{p}(N^{-1/2}T^{-1-\tau})=o_{p}(T^{-\tau}),$ (A.21)
@@
$\left\|\mathbf{I}\right\|=\left\|T^{-\tau}\left(\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right)\right\|=O_{p}(\xi_{N,T,\tau}^{-1})=o_{p}(T^{-\tau}),$ (A.22)
@@
$\left\|\mathbf{I}\mathbf{I}\right\|=\left\|T^{-\tau}\left(\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right)\right\|=O_{p}(\xi_{N,T,\tau}^{-1})=o_{p}(T^{-\tau}),$ (A.23)
@@
$\mathbf{III}$ $=T^{-\tau}\left(\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right)$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})\mathbf{X}_{i}$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}-\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}+\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}-\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{X}_{i}$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
@@
where we will start the term $\mathbf{III}_{d}$. Note that it explicitly looks like
$\mathbf{III}_{d}$ $=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$ (A.25)
with the key component $\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$, where
@@
as $T\to\infty$, which is positive definite constant matrix (see a similar result in A. 51 in the Supplement of Margaritella and Westerlund,, 2023). Therefore, clearly $\mathbf{III}_{d}$ converges to a positive semi-definite matrix, because it is a matrix-valued form. Importantly, it is not a zero matrix, because, according to Exercise 8.13 (b) in Abadir and Magnus, (2005), $\mathbf{III}_{d}=\mathbf{0}_{k\times k}$ is equivalent to $\mathrm{tr}(\mathbf{III}_{d})=0$ for a positive semi-definite matrix. Note that